# Phase error rate estimation with basis-efficiency mismatch for decoy-state BB84

### v1

@@ VII.2 Recipe for computing key rates in the presence of basis-efficiency mismatch
Thus, upper bounds on $\delta_{1},\delta_{2}$ can be computed using Eq. 78 and the bounds in Eq. 79. If one further assumes that $d_{\text{det}},\Delta_{\text{dc}},\Delta_{\eta}<<1$, the following simplified bounds can be obtained
$\delta_{1}$ $\lesssim 4\max\left\{\Delta_{\text{dc}},\Delta_{\eta}\right\},$ (80)
$\delta_{2}$ $\lesssim 2\max\left\{\Delta_{\text{dc}},\Delta_{\eta}\right\}.$
We are now finally ready to compute key rates. The key rate expression for the decoy-state BB84 protocol is given by Eq. 65. To use this expression, refer to Eqs. 51 and 48 (which are notationally equivalent). The bounds for the decoy analysis in Eq. 51 are in turn found in Eqs. 59, 60 and 61, whereas the bound for the phase error estimation is found in Eq. 44. For the value of $\delta_{1},\delta_{2}$, we use Eqs. 78 and 79. For the BB84 protocol where Alice sends single photons, the key rate is given by Eqs. 10 and 44.
VII.3 Plots
We plot finite size key rates for the decoy-state BB84 protocol described in Section VI.1. We choose typical protocol parameters and plot the key rate for the expected observations for a given channel model. For best results, one would optimize over the protocol parameter choices.
@@ VII.3 Plots
We set the basis choice probabilities to be $p^{(A)}_{(Z)}=p^{(B)}_{(Z)}=0.5$ and $p^{(A)}_{(X)}=p^{(B)}_{(X)}=0.5$, and $p_{\text{Z,T}}=0.05$ (probability of $Z$ basis rounds used for testing). We set the detector parameters to be $\eta_{\text{det}}=0.7$ and $d_{\text{det}}=10^{-6}$. We set the misalignment angle $\theta$ to be $2^{\circ}$. We set the number of bits used for error-correction to be $\lambda_{\text{EC}}(n_{X,\mu_{\vec{k}}},n_{K,\mu_{\vec{k}}},e^{\text{obs}}_{X,\mu_{\vec{k}}},e^{\text{obs}}_{Z})=f_{\text{EC}}n_{K}h(e^{\text{obs}}_{Z})$, where $f_{\text{EC}}=1.16$ is the error-correction efficiency. The decoy intensities are chosen to be $\mu_{1}=0.9$, $\mu_{2}=0.1$, and $\mu_{3}=0$. Each intensity is chosen with equal probability. We set $\varepsilon_{\text{AT-a}}=\varepsilon_{\text{AT-b}}=\varepsilon_{\text{AT-c}}=\varepsilon_{\text{AT-d}}=\varepsilon_{\text{EV}}=\varepsilon_{\text{PA}}=10^{-12}$. This leads to a value of $\varepsilon_{\text{AT}}=\sqrt{12}\times 10^{-12}$. The overall security parameter is then given by $(2\sqrt{12}+2)10^{-12}$. In Fig. 3, we plot the finite size key rate for various values of loss and detector characterizations $\Delta_{\eta},\Delta_{\text{dc}}$ for $n_{\text{total}}=10^{12}$ number of total signals.
Figure 3: Finite-size key rates in the presence of basis-efficiency mismatch, for the decoy-state BB84 protocol.
@@
For $\Delta_{\eta}=\Delta_{\text{dc}}=0\%$, we have $\delta_{1}=\delta_{2}=0$. Therefore the phase error rate bound from Eq. 44 reduces to the scenario where the basis-independent loss assumption is satisfied (Eq. 24). For non-zero values of $\Delta_{\eta},\Delta_{\text{dc}}$, the key rate is reduced. This is mostly due to the increase in the bound for the phase error rate from Eq. 44 from $\delta_{1}$, which is approximately $4\max\{\Delta_{\eta},\Delta_{\text{dc}}\}$ (Eq. 80).
VIII Detector Side-Channels
@@
where $d$ denotes the spatio-temporal mode, and $\Gamma_{\eta_{b_{i}}(d),d_{b_{i}}(d)}$ denotes the POVM element corresponding to that mode. The detector has loss $\eta_{b_{i}}(d)$ for this mode, and a dark count rate of $d_{b_{i}}(d)$.
The block-diagonal structure with respect to $d$ in the above equation reflects the fact that our model assumes no interference between any pair of spatio-temporal modes during the measurement process. Note that this assumption is physically motivated, and our proof can accommodate a more complicated model with off-diagonal blocks as well (as long as one can suitably bound $\delta_{1},\delta_{2}$). Due to the block-diagonal structure of the above POVM element Eq. 81, and the fact that $\delta_{1},\delta_{2}$ are $\infty$-norms which can be computed on each block-diagonal part separately, it is straightforward to see that our computation of $\delta_{1},\delta_{2}$ is directly applicable to the above scenario. In particular, the recipe from Section VII.2, and the computed key rates from Section VII.3 are valid for the scenario where Eve can control the value of $\eta_{b_{i}},d_{b_{i}}$ in the specified ranges (Eq. 74), via some extra spatio-temporal modes. Most importantly, our analysis does not depend on the number of such spatio-temporal modes. Thus, we are able to address scenarios where Eve has an arbitrary number of spatio-temporal modes, to induce (a bounded amount of) basis-efficiency mismatch in the detector.
@@
VIII.1 Independent and Correlated Detectors
@@
We first consider the case of independent detector behaviour in each round. In this case each round has a well-defined POVM that is independent of those in other rounds, i.e the POVM is tensor product with other rounds. Here we simply note that all the statistical claims used in our phase error estimation proofs (from Section B.2) remain true even if the POVM measurements are independent (but not IID). We comment on this and restate some of our lemmas for independent (but not identical) measurements in Section B.3. Moreover the Serfling statement (Lemma 2) does not assume any IID property of the input string. Thus, our bounds on the phase error rate remain unchanged as long as $\delta_{1},\delta_{2}$ can be bounded for the independent POVMs in each round. This deals with the case of independent detectors.
@@
IX Summary and Discussion
In this work, we presented a finite-size security proof of QKD in the presence of imperfectly characterized and adversary controlled basis-efficiency mismatch. Thus, we addressed a longstanding assumption made in security proofs within the EUR and phase error correction frameworks. Since our methods permit adversarial control over the efficiency mismatch, we also resolved a critical open problem of detector side-channels, which has remained unresolved in existing security proof approaches for standard QKD. Furthermore, we rigorously establish the security for variable-length QKD protocols within the EUR framework, which has significant practical consequences. Finally, we also fixed several technical issues in the security analysis of decoy-state QKD within the EUR framework. We applied our results to the decoy-state BB84 protocol, demonstrating practical key rates in the finite-size regime even in the presence of basis-efficiency mismatch. Taken together, these results are a significant step towards protocol security of the EUR proof technique, and the implementation security of QKD using trusted detection setups.
There are several opportunities for interesting work that builds upon these results. One possibility is to integrate our methods with established methods for addressing source imperfections [6, 7, 8]. Such a combination would lead to a security proof robust to both source and detector imperfections. Another intriguing avenue is applying these methods to passive detection setups, where the basis-efficiency mismatch assumption translates to an assumption of a perfectly balanced beam splitter and identical detectors.
For computing key rates based on our results, suitable bounds on $\delta_{1},\delta_{2}$ are required. While our current analysis relies on some simplifications to obtain these bounds, they can likely be improved. Additionally, our analysis shows anomalous behaviour in specific regimes. For instance, when the minimum dark count rate is set to zero, then our methods, as presented, break down and yield zero key rate (because $\Delta_{\text{dc}}=1$). We expect a more refined analysis to be able to get rid of such oddities. Finally, an ambitious direction of future work would be to directly address correlated detectors. While our methods probably allow for one way of addressing this (Section VIII.1), we find this unsatisfactory due to the requirement of protocol modifications.
@@ Appendix A Technical Statements
We use $S_{\circ}(Q)$ to denote the set of normalized states on $Q$. We use $S_{\bullet}(Q)$ to denote the set of all sub-normalized states on $Q$. We use $\text{Pos}(Q)$ to denote the set of positive semi-definite operators on $Q$. The smoothing on the min and max entropies is with respect to the purified distance [35, Definition 3.8].
@@
Appendix B Sampling

### v2

@@ VI.3.1 Active BB84 detection setup without any hardware modification
Thus, upper bounds on $\delta_{1},\delta_{2}$ can be computed using Eq. 60 and the bounds in Eq. 61. It is these bounds that we use to compute key rates.
VI.3.2 Random Swapping of 0 and 1 Detectors
In [12] it was argued that random swapping of the 0 and the 1 detector can be used to remove basis-efficiency mismatch for single-photon pulses entering Bob’s detectors. Note that this trick only works for the single-photon subspace. We will now adapt our analysis to the case where Bob randomly swaps the 00 and the $1$ detector.
In the scenario where we randomly swap the 00 and the $1$ detectors, we make certain physically motivated assumptions (Eq. 62) about the detector setup. In particular, we assume that the dark count rate is a property of the detector only. Furthermore, we assume that the basis choice setting does not change the detector parameters. This means that the dark count rate and detection efficiency in both bases is the same (though these can be different for each detector). Thus, we have
$\eta_{X_{0}}$ $=\eta_{Z_{0}}\eqqcolon\eta_{0}$ (62)
$\eta_{X_{1}}$ $=\eta_{Z_{1}}\eqqcolon\eta_{1}$
$d_{X_{0}}$ $=d_{Z_{0}}\eqqcolon d_{0}$
$d_{X_{1}}$ $=d_{Z_{1}}\eqqcolon d_{1}.$
We will see that this indeed allows us to obtain improved results, even though it does not completely remove efficiency mismatch. In particular, the leading order terms in $\delta_{1},\delta_{2}$ are improved in the new bounds obtained in Eqs. 64 and 65. Note that our metrics $\delta_{1},\delta_{2}$ do not improve unless we make these assumptions. These assumptions are also implicit in the claims presented in Ref. [12].
If the random swapping is implemented with probability $p$, the Bob’s POVM elements are given by
$\Gamma^{(B),(\text{swap})}_{(b,\bot)}$ $=\sum_{N_{0},N_{1}=0}^{\infty}(1-d_{b_{0}})(1-d_{b_{1}})\left((1-p)(1-\eta_{b_{0}})^{N_{0}}(1-\eta_{b_{1}})^{N_{1}}+p(1-\eta_{b_{1}})^{N_{0}}(1-\eta_{b_{0}})^{N_{1}}\right)\outerproduct{N_{0},N_{1}}{N_{0},N_{1}}_{b}$ (63)
$\Gamma^{(B),(\text{swap})}_{(b,0)}$ $=\Big{(}(1-p)(1-d_{b_{1}})\sum_{N_{0},N_{1}=0}^{\infty}(1-(1-d_{b_{0}})(1-\eta_{b_{0}})^{N_{0}})(1-\eta_{b_{1}})^{N_{1}}$
$+p(1-d_{b_{0}})\sum_{N_{0},N_{1}=0}^{\infty}(1-(1-d_{b_{1}})(1-\eta_{b_{1}})^{N_{0}})(1-\eta_{b_{0}})^{N_{1}}\Big{)}\outerproduct{N_{0},N_{1}}{N_{0},N_{1}}_{b}+\frac{1}{2}\Gamma^{(B)}_{(b,\text{dc})}$
$\Gamma^{(B),(\text{swap})}_{(b,1)}$ $=\Big{(}(1-p)(1-d_{b_{0}})\sum_{N_{0},N_{1}=0}^{\infty}(1-\eta_{b_{0}})^{N_{0}}(1-(1-d_{b_{1}})(1-\eta_{b_{1}})^{N_{1}})$
$+p(1-d_{b_{1}})\sum_{N_{0},N_{1}=0}^{\infty}(1-\eta_{b_{1}})^{N_{0}}(1-(1-d_{b_{0}})(1-\eta_{b_{0}})^{N_{1}})\Big{)}\outerproduct{N_{0},N_{1}}{N_{0},N_{1}}_{b}+\frac{1}{2}\Gamma^{(B)}_{(b,\text{dc})},$
analogously to Eq. 57. Alice and Bob’s joint POVM elements can be constructed from Eqs. 2, 3 and 63 analogously to Eq. 58. Therefore we can repeat the calculations for $\delta_{1},\delta_{2}$ using the recipe from Section VI.1. We explain these computations in Appendix H and obtain (for swap probability $p=1/2$)
$\delta_{1}$ $\leq 4\left(1-\sqrt{1-(1-d_{\text{mult}})^{2}\frac{(1-r_{\eta})^{2}}{2}}\right),$ (64)
$\delta_{2}$ $\leq(1-d_{\text{mult}})^{2}\frac{(1-r_{\eta})^{2}}{2},$
where
$d_{\text{mult}}$ $=1-\sqrt{(1-d_{0})(1-d_{1})}\geq d_{\text{min}},$ (65)
$r_{\eta}$ $=\frac{\eta_{\text{min}}}{\eta_{\text{max}}}\geq\frac{1-\Delta_{\eta}}{1+\Delta_{\eta}}.$
Thus, upper bounds on $\delta_{1},\delta_{2}$ in case of random swapping of detectors can be computed using Eq. 64 and the bounds in Eq. 65. We see that these bounds are better than the earlier bounds from Eq. 60. On inspecting our calculations from Appendix H, we find that the zero-photon component of $\delta_{1},\delta_{2}$ goes to zero due to $d_{X}=d_{Z}$. Furthermore, random swapping in addition to the assumption of $\eta_{X}=\eta_{Z}$ leads to the single-photon contribution also being zero. Thus, we are left with the two-photon contribution.
VI.4 Plots
Figure 4: Finite-size key rates in the presence of basis-efficiency mismatch, for the decoy-state BB84 protocol, against loss. We plot key rates for $N_{\text{tot}}=10^{12}$ number of total signals sent, for various values of $\Delta_{\eta},\Delta_{\text{dc}}$. We find that random swapping of the 00 and $1$ detectors drastically improves the key rates obtained.
@@ VI.4 Plots
We plot finite size key rates for the decoy-state BB84 protocol described in Section V.1. We choose typical protocol parameters and plot the key rate for the expected observations for a given channel model. For best results, one would optimize over the protocol parameter choices. For all plots, we set the basis choice probabilities to be $p^{(A)}_{(Z)}=p^{(B)}_{(Z)}=0.5$ and $p^{(A)}_{(X)}=p^{(B)}_{(X)}=0.5$, and $p_{\text{Z,T}}=0.05$ (probability of $Z$ basis rounds used for testing). We set the detector parameters to be $\eta_{\text{det}}=0.7$ and $d_{\text{det}}=10^{-6}$. We set the misalignment angle $\theta$ to be $2^{\circ}$. We set the number of bits used for error-correction to be $\lambda_{\text{EC}}(n_{X,\mu_{\vec{k}}},n_{K,\mu_{\vec{k}}},e^{\text{obs}}_{X,\mu_{\vec{k}}},e^{\text{obs}}_{Z})=f_{\text{EC}}n_{K}h(e^{\text{obs}}_{Z})$, where $f_{\text{EC}}=1.16$ is the error-correction efficiency. The decoy intensities are chosen to be $\mu_{1}=0.9$, $\mu_{2}=0.1$, and $\mu_{3}=0$. Each intensity is chosen with equal probability. We set $\varepsilon_{\text{AT-a}}=\varepsilon_{\text{AT-b}}=\varepsilon_{\text{AT-c}}=\varepsilon_{\text{AT-d}}=\varepsilon_{\text{EV}}=\varepsilon_{\text{PA}}=10^{-12}$. This leads to a value of $\varepsilon_{\text{AT}}=\sqrt{12}\times 10^{-12}$. The overall security parameter is then given by $(2\sqrt{12}+2)10^{-12}$. Due to machine precision issues arising from small values of $\varepsilon_{\text{AT}}^{2}$, we use Hoeffdings inequality to bound $\gamma_{\text{bin}}$ (Eq. 27) instead of using the cumulative binomial distribution (which is tighter).
  1. 1.
@@
    In Fig. 4, we plot the finite size key rate against loss for various values of detector characterizations $\Delta_{\eta},\Delta_{\text{dc}}$ for $n_{\text{total}}=10^{12}$ number of total signals. For $\Delta_{\eta}=\Delta_{\text{dc}}=0$, we have $\delta_{1}=\delta_{2}=0$. Therefore the phase error rate bound from Eq. 34 reduces to the scenario where the basis-independent loss assumption is satisfied (Eq. 14). For non-zero values of $\Delta_{\eta},\Delta_{\text{dc}}$, the key rate is reduced. This is mostly due to the increase in the bound for the phase error rate from Eq. 34 from $\delta_{1}$. We find that random swapping leads to a dramatic improvement in performance.
  2. 2.
    In Fig. 5, we plot the finite size key rate against loss for various values of total signals sent. We set $\Delta_{\eta}=\Delta_{\text{dc}}=0.05$. We find that we get close to asymptotic key rates already at $N_{\text{tot}}=10^{12}$ signals sent.
  3. 3.
    In Fig. 6, we plot the finite size key rate against detector characterization parameters $\Delta_{\eta},\Delta_{\text{dc}}$. We find that our methods can tolerate a significant amount of error in detector characterization. In fact, with random swapping of detectors, we get positive key rate for $N_{\text{tot}}=10^{12}$ signals sent for $\Delta_{\eta},\Delta_{\text{dc}}$ upto $0.35$.
Figure 5: Finite-size key rates in the presence of basis-efficiency mismatch, for the decoy-state BB84 protocol against loss. We plot key rates for various values of total number of signals sent ($N_{\text{tot}}$), for $\Delta_{\eta}=\Delta_{\text{dc}}=0.05$.
Figure 6: Finite-size key rates in the presence of basis-efficiency mismatch, for the decoy-state BB84 protocol against detector characterization parameters $\Delta_{\eta},\Delta_{\text{dc}}$. We plot key rates for a channel with $25$dB loss.
VII Detector Side-Channels
@@
where $d$ denotes the spatio-temporal mode, and $\Gamma_{\eta_{b_{i}}(d),d_{b_{i}}(d)}$ denotes the POVM element corresponding to that mode. The detector has loss $\eta_{b_{i}}(d)$ for this mode, and a dark count rate of $d_{b_{i}}(d)$. The block-diagonal structure with respect to $d$ in the above equation reflects the fact that our model assumes no interference between any pair of spatio-temporal modes during the measurement process. Alternatively, one can equivalently consider a model where $d$ in Eq. 66 is interpreted as a classical flag, which Eve can use to choose detector efficiencies and dark count rates in their allowed ranges. While we only consider models of the above form in this work, our proof can accommodate a more complicated model with off-diagonal blocks as well (as long as one can suitably bound $\delta_{1},\delta_{2}$).
Due to the block-diagonal structure of the above POVM element Eq. 66, and the fact that $\delta_{1},\delta_{2}$ are $\infty$-norms which can be computed on each block-diagonal part separately, it is straightforward to see that our computation of $\delta_{1},\delta_{2}$ is directly applicable to the above scenario. In particular, the recipe from Section VI.1, and the computed key rates from Section VI.4 are valid for the scenario where Eve can control the value of $\eta_{b_{i}},d_{b_{i}}$ in the specified ranges (Eq. 59), via some extra spatio-temporal modes. Most importantly, our analysis does not depend on the number of such spatio-temporal modes. Thus, we are able to address scenarios where Eve has an arbitrary number of spatio-temporal modes, to induce (a bounded amount of) basis-efficiency mismatch in the detector.
@@
VII.1 Independent and Correlated Detectors
@@
We first consider the case of independent detector behaviour in each round. In this case each round has a well-defined POVM that is independent of those in other rounds, i.e the POVM is tensor product with other rounds. Here we simply note that all the statistical claims used in our phase error estimation proofs (from Section C.2) remain true even if the POVM measurements are independent (but not IID). We comment on this and restate some of our lemmas for independent (but not identical) measurements in Section C.3. Moreover the Serfling statement (Lemma 2) does not assume any IID property of the input string. Thus, our bounds on the phase error rate remain unchanged as long as $\delta_{1},\delta_{2}$ can be bounded for the independent POVMs in each round. This deals with the case of independent detectors.
@@
VIII Summary and Discussion
In this work, we presented a finite-size security proof of QKD in the presence of imperfectly characterized and adversary controlled basis-efficiency mismatch. Thus, we addressed a longstanding assumption made in security proofs within the EUR and phase error correction frameworks. Before this work, proofs within these frameworks were not stable, and would be invalidated by infinitesimal amounts of basis-efficiency mismatch. Since our methods permit adversarial control over the efficiency mismatch, we also resolved a critical open problem of detector side-channels, which has remained unresolved in existing security proof approaches for standard QKD. Finally, we also fixed several technical issues in the security analysis of decoy-state QKD within the EUR framework.
We applied our results to the decoy-state BB84 protocol, demonstrating practical key rates in the finite-size regime even in the presence of basis-efficiency mismatch. We also investigate quantitatively, the effect of methods such as random swapping of detectors, to reduce efficiency mismatch. Taken together, these results are a significant step towards protocol security of the EUR proof technique, and the implementation security of QKD using trusted detection setups.
There are several opportunities for interesting work that builds upon these results. One possibility is to integrate our methods with established methods for addressing source imperfections [6, 7, 8]. Such a combination would lead to a security proof robust to both source and detector imperfections. Another avenue is applying these methods to passive detection setups, where the basis-efficiency mismatch assumption translates to an assumption of a perfectly balanced beam splitter and identical detectors. For computing key rates based on our results, suitable bounds on $\delta_{1},\delta_{2}$ are required. While our current analysis relies on some simplifications to obtain these bounds, they can likely be improved. Finally, an ambitious direction of future work would be to directly address correlated detectors. While our methods probably allow for one way of addressing this (Section VII.1), we find this unsatisfactory due to the requirement of protocol modifications.
@@ Appendix A Technical Statements
We use $S_{\circ}(Q)$ to denote the set of normalized states on $Q$. We use $S_{\bullet}(Q)$ to denote the set of all sub-normalized states on $Q$. We use $\text{Pos}(Q)$ to denote the set of positive semi-definite operators on $Q$. The smoothing on the min and max entropies is with respect to the purified distance [54, Definition 3.8].
@@
Appendix B Variable-length security proof
In this appendix we will prove the following theorem regarding variable-length security of the protocol from Section II.
See 1
Proof.
Our proof will consist of three parts. In the first part, we will discuss the security definition for variable-length QKD protocols. In the second part we will use entropic uncertainty relations and Eq. 5 to obtain a suitable lower bound on the smooth min entropy of the raw key register in the QKD protocol. In the third part, we use this to prove variable-length security.
B.0.1 Variable-length security definition
In order to prove the $\varepsilon_{\text{secure}}$-security of a variable-length QKD protocol [23], one must show that for all attacks by the adversary, the following statement is true :
B.0.2 Bounding the smooth min entropy
Let us turn our attention to the phase error rate estimate. Let
$\leq\varepsilon_{\text{AT}}^{2},$
where the sum is over all possible values of $n_{X},n_{K},e^{\text{obs}}_{X},e^{\text{obs}}_{Z}$. We will utilize Eq. 74, which follows from Eq. 5, in bounding the smooth min entropy of the raw key register.
To do so, focus on the state $\rho_{A^{n_{K}}B^{n_{K}}E^{n}C^{n}|\Omega_{(n_{X},n_{K},e^{\text{obs}}_{X},e^{\text{obs}}_{Z})}}$, which is the state on the detected key generation rounds. This state can be obtained by transforming Bob’s measurement procedure to consist of two steps, and then only implementing the first step measurement which determines the detect vs no-detect outcome. Such a state can be rigorously obtained using Lemma 1 from Section III. For the purposes of this proof, we only need the fact that it is well defined. We will obtain a bound on the smooth min entropy of the key generated from this state.
Remark 15.
We can make Bob measure his systems $B$ in the $X$ basis to obtain the classical outcome $X_{B}$. Then, using data processing [54, Theorem 6.2], we obtain
which along with Eq. 76 gives us
B.0.3 Proving variable-length security
$\leq\varepsilon_{\text{PA}}+2\varepsilon_{\text{AT}}.$
Here, we used the leftover-hashing lemma [2, Proposition 9] with the appropriate smoothing parameter on the sub-normalized state $(\rho_{|\Omega_{(n_{X},n_{K},e^{\text{obs}}_{X},e^{\text{obs}}_{Z})}})_{\wedge\Omega_{\text{EV}}}$ for the first inequality. Since we use sub-normalized conditioning on $\Omega_{\text{EV}}$, it only appears in the smooth min entropy term and not inside the probability. Next, we use [2, Lemma 10] to get rid of the sub-normalized conditioning ($\wedge\Omega_{\text{EV}}$) in the smooth min entropy term in the second inequality. We used [54, Lemma 6.8] to split off the error-correction information ($\lambda_{\text{EC}}(n_{X},n_{K},e^{\text{obs}}_{X},e^{\text{obs}}_{Z})$) and error-verification information ($\log(2/\varepsilon_{\text{EV}})$) in the third inequality. We used the bound on the smooth min entropy from Eq. 78 for the fourth inequality, and the values of $l(n_{X},n_{K},e^{\text{obs}}_{X},e^{\text{obs}}_{Z})$ and $\lambda_{\text{EC}}(n_{X},n_{K},e^{\text{obs}}_{X},e^{\text{obs}}_{Z})$ from Eq. 6 for the fifth equality. We used concavity of the square root function and Jensen’s inequality to pull the sum over probabilities inside the square root for the sixth inequality, and Eq. 74 for the final inequality.
Remark 16.
Note that the critical step here was using the concavity of the square root function to see that a bound on the average failure probability of the phase error estimation procedure is enough to prove security. This is the same fundamental trick used by Ref. [27, 28, 8]. Our presentation here is in the EUR framework, where this manifests in the deliberate choice of our smoothing parameter in the first part of the proof.
Remark 17.
It is interesting to note that these issues are completely avoided by the above proof, in a very different manner than [24]. This is because in this proof, we always apply the leftover-hashing lemma on a state conditioned on the specific length of the raw key register (Eq. 78). Therefore, the leftover-hashing lemma can be applied in a straightforward manner, and there are no issues is choosing the hashing family based on the specific length of the raw key register. In other words, the PA procedure described above is valid for this proof. In a similar sense, the variable-length security proof from [24] critically relied on a technical lemma (Lemma 9), that necessitated the use of Rényi entropies instead of smooth min entropy in that work. However, the variable-length proof presented above takes a different approach, and does not impose the same requirements on the behaviour of smooth min entropy. Again, this is due to the use of Eq. 78.
B.1 Comparing to fixed-length protocols
Appendix C Sampling